Differentiate:



f(x)= (2x⁴+5x+2)



g(x)= 3x²√6x³+5x²+1



h(x)= 4x²/√x+7

1
Expert's answer
2022-05-30T23:29:04-0400

a)


"f'(x)= 2(4x^3)+5(1)+0"

"=8x^3+5"

b)


"g'(x)=(3x^2)'\sqrt{6x^3+5x^2+1}"

"+3x^2(\sqrt{6x^3+5x^2+1})'"

"=6x\sqrt{6x^3+5x^2+1}"

"+\dfrac{3x^2}{2\sqrt{6x^3+5x^2+1}}(6x^3+5x^2+1)'"


"=6x\sqrt{6x^3+5x^2+1}"

"+\dfrac{3x^2(18x^2+10x)}{2\sqrt{6x^3+5x^2+1}}"

"=\dfrac{3x(12x^3+10x^2+2+9x^3+5x^2)}{\sqrt{6x^3+5x^2+1}}"

"=\dfrac{3x(21x^3+15x^2+2)}{\sqrt{6x^3+5x^2+1}}"

c)


"h'(x)=\dfrac{(4x^2)'\sqrt{x+7}-4x^2(\sqrt{x+7})'}{x+7}"

"=\dfrac{8x\sqrt{x+7}-4x^2(\dfrac{1}{2\sqrt{x+7}})}{x+7}"

"=\dfrac{8x^2+56x-2x^2}{(x+7)^{3/2}}"

"=\dfrac{6x^2+56x}{(x+7)^{3/2}}"